# An extension of Viennot’s shadow to rook placements via orbit harmonics

### v3

@@ 1. Introduction
The method of orbit harmonics dates back to Kostant  [6]. It has seen its application in numerous settings, including cohomology theory  [3], Macdonald theory  [4, 5], cyclic sieving [10], Donaldson–Thomas invariants  [11], and Ehrhart theory  [12]. In cases where the locus $\mathcal{Z}$} has nice combinatorial structures or symmetries, one can usually expect that the algebraic properties of $R({\mathcal{Z}})$ are governed by the combinatorial properties of $\mathcal{Z}$}.
Let ${\mathbf{x}}_{n\times m}=(x_{i,j})_{1\leq i\leq n,\,1\leq j\leq m}$ be an $n\times m$ matrix of variables. Let ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ be the polynomial ring over these variables and ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices. The application of orbit harmonics to finite matrix loci ${\mathcal{Z}}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ was initiated by Rhoades [13]. He considered the case $n=m$ with the point locus ${\mathcal{Z}}={\mathfrak{S}}_{n}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of $n\times n$ permutation matrices. This permutation matrix locus carries an action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$ by left and right multiplication. Algebraic properties of $R({\mathfrak{S}}_{n})$ are governed by the longest increasing subsequences in ${\mathfrak{S}}_{n}$ and Viennot’s shadow line construction. Liu [7] extended this work to the locus ${\mathcal{Z}}={\mathbb{Z}}_{r}\wr{\mathfrak{S}}_{n}$ of $r$-colored permutation matrices. Liu, Ma, Rhoades, Zhu  [8] further studied the matrix loci ${\mathcal{Z}}=\{w\in{\mathfrak{S}}_{n}:w^{2}=1\}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of involutions and the matrix loci ${\mathcal{Z}}=\mathcal{PM}_{n}=\{w\in{\mathcal{M}}_{n}:w(i)\neq i\text{ for all }i\}$ of fixed-point-free involutions.
@@ 2.1. Gröbner Theory
forms a basis for the vector space ${\mathbb{C}}[{\mathbf{x}}_{N}]/I$. This is called the standard monomial basis. See [2] for further details on Gröbner theory.
@@ 2.3. Schensted Correspondence
Denote by $\mathrm{SYT}(\lambda)$ the set of standard Young tableaux of shape $\lambda$. The Schensted correspondence [15] is a bijection
@@
that takes a permutation in ${\mathfrak{S}}_{n}$ to a pair of standard Young tableaux of the same shape. This bijection is usually achieved through an insertion algorithm, details of which can be found in  [14]. In Section 4.1, we will introduce an alternate description of the Schensted correspondence discovered by Viennot [16].
@@ 2.4. Representation Theory
The degree-$n$ component $\Lambda_{n}$ of $\Lambda$ has bases indexed by partitions $\lambda\vdash n$. For example, the complete homogeneous basis of $\Lambda_{n}$ is $\{h_{\lambda}\,:\,\lambda\vdash n\}$ where $h_{\lambda}=h_{\lambda_{1}}h_{\lambda_{2}}\dots$ for $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$. We will also need the Schur basis $\{s_{\lambda}\,:\,\lambda\vdash n\}$ (see [9] for its definition). If $\lambda=(n)$ one has $s_{(n)}=h_{n}$.
@@
Thus, $\eta_{j}\cdot V^{\lambda}\neq 0$ if and only if $(\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda})^{{\mathfrak{S}}_{j}}\neq 0$. This is true if and only if the trivial representation $V^{(j)}$ appears with positive multiplicity in the decomposition of $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$. The branching rule of ${\mathfrak{S}}_{n}$-modules (see [9, 14]) states that
@@ 3.1. $r=0$ case
When $r=0$, write ${\mathcal{Z}}_{n,m}:={\mathcal{UZ}}_{n,m,0}$ for the locus of all rook placements on the $n\times m$ board. The locus ${\mathcal{Z}}_{n,m}$ is a finite shifted locus, which was defined and studied by Reiner and Rhoades [12, Sec. 3.4]. Lemma 3.3 and Proposition  3.4 below can be derived directly from the results by Reiner and Rhoades, but we still include their proofs as they will promote understanding of later sections.
@@ 4.1. Viennot Shadow
We begin this section by describing Viennot’s shadow line construction [16]. Let $w=[w(1),w(2),\dots,w(n)]\in{\mathfrak{S}}_{n}$ be a permutation. We draw its diagram on an $n\times n$ grid by putting a point at $(i,w(i))$ for $i\in[n]$. For example, the permutation $w=[6,3,5,7,1,2,8,4]\in{\mathfrak{S}}_{8}$ has the digram shown below.
@@
From the construction, it is immediate that the shadow set ${\mathcal{S}}(w)$ of any permutation $w$ is a rook placement. However, not all rook placements are shadow sets of permutations. In  [13], Rhoades described the following algorithm to determine whether a rook placement is the shadow set of some permutation. In the next section, we present a generalized version of this algorithm.
@@ 4.2. Extended shadow sets
Let $(P,Q)$ be the pair of tableaux corresponding to ${\mathcal{EX}}({\mathcal{R}})$ in the Schensted correspondence. We adopt the notion in Remark 4.3 so the content of $P$ is $\{\overline{1},\overline{2},\dots,\overline{n-r},1,2,\dots m\}$ and the content of $Q$ is $\{\overline{1},\overline{2},\dots,\overline{m-1},1,2,\dots n\}$. According to Viennot’s construction in  [16], we have that $\overline{1},\overline{2},\dots,\overline{n-r}$ are in the first row of $P$ as they are the coordinates of the infinite horizontal rays of shadow lines $L_{1},L_{2},\dots,L_{n-r}$. Similarly, $\overline{1},\overline{2},\dots,\overline{m-r}$ are in the first row of $Q$. As the horizontal coordinates of points in the shadow set are entries in $P$ beyond the first row, and the vertical coordinates of points in the shadow set are entries in $Q$ beyond the shadow set, we conclude that ${\mathcal{ES}}({\mathcal{R}})={\mathcal{S}}({\mathcal{EX}}({\mathcal{R}}))\subseteq[n]\times[m]$. ∎
@@ 4.4. Standard monomial basis and Hilbert series
Let $k$ be the number of shadow lines of $w$ that we use to construct ${\mathcal{S}}(w)$. Write $w\mapsto(P,Q)$ according to the Schensted correspondence, where $P,Q$ have the same shape $\lambda$. The shadow line construction in Subsection 4.1 implies that $k=\lambda_{1}$. In addition, Schensted correspondence possesses the standard property [15, Thm. 1] $\lambda_{1}=\mathrm{lis}(w)$, so we have $k=\mathrm{lis}(w)$ and thus Equation (4.4) holds. ∎
@@ 6. Conclusion
We provide an open problem regarding log-concavity. Recall that a sequence of positive real numbers $(a_{1},a_{2},\dots,a_{N})$ is log-concave if for all $1<i<N$ we have $a_{i}^{2}\geq a_{i-1}\cdot a_{i+1}$. Chen conjectured [1] that the sequence $(a_{n,1},a_{n,2},\dots,a_{n,n})$ is log-concave where $a_{n,k}$ is the number of permutations in ${\mathfrak{S}}_{n}$ with longest subsequence of length $k$. Using the terminology of orbit harmonics, Chen’s conjecture means that the Hilbert series of the orbit harmonics ring $R({\mathfrak{S}}_{n})$ of the permutation matrix locus ${\mathfrak{S}}_{n}$ is log-concave.
We can generalize the definition of log-concavity by incorporating group action. Let $G$ be a group and $(V_{1},V_{2},\dots,V_{N})$ is a sequence of $G$-modules. We say that $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave if for all $1<i<N$ we have a $G$-module surjection $V_{i}\otimes V_{i}\twoheadrightarrow V_{i-1}\otimes V_{i+1}$ where the $G$-action is diagonal, i.e., $g\cdot(v\otimes w)=(g\cdot v)\otimes(g\cdot w)$. A graded $G$-module $V=\bigoplus_{d=0}^{N}V_{d}$ is $G$-equivariant log-concave if the sequence $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave. Rhoades conjectured [13] that $R({\mathfrak{S}}_{n})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-equivariant log-concave, which generalizes Chen’s conjecture.
@@
Another interesting future direction is to study the orbit harmonics ring $R({\mathcal{Z}}_{n,m,r})$. Since ${\mathcal{Z}}_{n,m,r}\subseteq{\mathcal{UZ}}_{n,m,r}$, we need to add some elements to the generating set in Definition 3.1 to obtain the defining ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}}_{n,m,r})$, such as $\sum_{i=1}^{n}\sum_{j=1}^{m}x_{i,j}$ and ${\mathfrak{m}}({\mathcal{R}})$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r+1}$. Some relevant results are presented in Zhu [17] focusing on the graded module structure.

### v4

@@ 1. Introduction
The method of orbit harmonics dates back to Kostant  [Kostant1963LieGR]. It has seen its application in numerous settings, including cohomology theory  [GARSIA199282], Macdonald theory  [griffin2021orderedsetpartitionsgarsiaprocesi, HAGLUND2018851], cyclic sieving [Oh-Rhoades], Donaldson–Thomas invariants  [Reineke-Rhoades], and Ehrhart theory  [reiner2024harmonicsgradedehrharttheory]. In cases where the locus $\mathcal{Z}$} has nice combinatorial structures or symmetries, one can usually expect that the algebraic properties of $R({\mathcal{Z}})$ are governed by the combinatorial properties of $\mathcal{Z}$}.
Let ${\mathbf{x}}_{n\times m}=(x_{i,j})_{1\leq i\leq n,\,1\leq j\leq m}$ be an $n\times m$ matrix of variables. Let ${\mathbb{C}}[{\mathbf{x}}_{n\times m}]$ be the polynomial ring over these variables and ${\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ be the affine space of $n\times m$ complex matrices. The application of orbit harmonics to finite matrix loci ${\mathcal{Z}}\subseteq{\mathrm{Mat}}_{n\times m}({\mathbb{C}})$ was initiated by Rhoades [rhoades2024increasing]. He considered the case $n=m$ with the point locus ${\mathcal{Z}}={\mathfrak{S}}_{n}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of $n\times n$ permutation matrices. This permutation matrix locus carries an action of ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$ by left and right multiplication. Algebraic properties of $R({\mathfrak{S}}_{n})$ are governed by the longest increasing subsequences in ${\mathfrak{S}}_{n}$ and Viennot’s shadow line construction. Liu [Liu] extended this work to the locus ${\mathcal{Z}}={\mathbb{Z}}_{r}\wr{\mathfrak{S}}_{n}$ of $r$-colored permutation matrices. Liu, Ma, Rhoades, Zhu  [Liu-Ma-Rhoades-Zhu] further studied the matrix loci ${\mathcal{Z}}=\{w\in{\mathfrak{S}}_{n}:w^{2}=1\}\subseteq{\mathrm{Mat}}_{n\times n}({\mathbb{C}})$ of involutions and the matrix loci ${\mathcal{Z}}=\mathcal{PM}_{n}=\{w\in{\mathcal{M}}_{n}:w(i)\neq i\text{ for all }i\}$ of fixed-point-free involutions.
@@ 2.1. Gröbner Theory
forms a basis for the vector space ${\mathbb{C}}[{\mathbf{x}}_{N}]/I$. This is called the standard monomial basis. See [Cox-Little-O'shea] for further details on Gröbner theory.
@@ 2.3. Schensted Correspondence
Denote by $\mathrm{SYT}(\lambda)$ the set of standard Young tableaux of shape $\lambda$. The Schensted correspondence [Schensted_1961] is a bijection
@@
that takes a permutation in ${\mathfrak{S}}_{n}$ to a pair of standard Young tableaux of the same shape. This bijection is usually achieved through an insertion algorithm, details of which can be found in  [sagan2013symmetric]. In Section 4.1, we will introduce an alternate description of the Schensted correspondence discovered by Viennot [Viennot].
@@ 2.4. Representation Theory
The degree-$n$ component $\Lambda_{n}$ of $\Lambda$ has bases indexed by partitions $\lambda\vdash n$. For example, the complete homogeneous basis of $\Lambda_{n}$ is $\{h_{\lambda}\,:\,\lambda\vdash n\}$ where $h_{\lambda}=h_{\lambda_{1}}h_{\lambda_{2}}\dots$ for $\lambda=(\lambda_{1}\geq\lambda_{2}\geq\dots)$. We will also need the Schur basis $\{s_{\lambda}\,:\,\lambda\vdash n\}$ (see [Macdonald] for its definition). If $\lambda=(n)$ one has $s_{(n)}=h_{n}$.
@@
Thus, $\eta_{j}\cdot V^{\lambda}\neq 0$ if and only if $(\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda})^{{\mathfrak{S}}_{j}}\neq 0$. This is true if and only if the trivial representation $V^{(j)}$ appears with positive multiplicity in the decomposition of $\mathrm{Res}^{{\mathfrak{S}}_{n}}_{{\mathfrak{S}}_{j}}\,V^{\lambda}$. The branching rule of ${\mathfrak{S}}_{n}$-modules (see [Macdonald, sagan2013symmetric]) states that
@@ 3.1. $r=0$ case
When $r=0$, write ${\mathcal{Z}}_{n,m}:={\mathcal{UZ}}_{n,m,0}$ for the locus of all rook placements on the $n\times m$ board. The locus ${\mathcal{Z}}_{n,m}$ is a finite shifted locus, which was defined and studied by Reiner and Rhoades [reiner2024harmonicsgradedehrharttheory, Sec. 3.4]. Lemma 3.3 and Proposition  3.4 below can be derived directly from the results by Reiner and Rhoades, but we still include their proofs as they will promote understanding of later sections.
@@ 4.1. Viennot Shadow
We begin this section by describing Viennot’s shadow line construction [Viennot]. Let $w=[w(1),w(2),\dots,w(n)]\in{\mathfrak{S}}_{n}$ be a permutation. We draw its diagram on an $n\times n$ grid by putting a point at $(i,w(i))$ for $i\in[n]$. For example, the permutation $w=[6,3,5,7,1,2,8,4]\in{\mathfrak{S}}_{8}$ has the digram shown below.
@@
From the construction, it is immediate that the shadow set ${\mathcal{S}}(w)$ of any permutation $w$ is a rook placement. However, not all rook placements are shadow sets of permutations. In  [rhoades2024increasing], Rhoades described the following algorithm to determine whether a rook placement is the shadow set of some permutation. In the next section, we present a generalized version of this algorithm.
@@ 4.2. Extended shadow sets
Let $(P,Q)$ be the pair of tableaux corresponding to ${\mathcal{EX}}({\mathcal{R}})$ in the Schensted correspondence. We adopt the notion in Remark 4.3 so the content of $P$ is $\{\overline{1},\overline{2},\dots,\overline{n-r},1,2,\dots m\}$ and the content of $Q$ is $\{\overline{1},\overline{2},\dots,\overline{m-1},1,2,\dots n\}$. According to Viennot’s construction in  [Viennot], we have that $\overline{1},\overline{2},\dots,\overline{n-r}$ are in the first row of $P$ as they are the coordinates of the infinite horizontal rays of shadow lines $L_{1},L_{2},\dots,L_{n-r}$. Similarly, $\overline{1},\overline{2},\dots,\overline{m-r}$ are in the first row of $Q$. As the horizontal coordinates of points in the shadow set are entries in $P$ beyond the first row, and the vertical coordinates of points in the shadow set are entries in $Q$ beyond the shadow set, we conclude that ${\mathcal{ES}}({\mathcal{R}})={\mathcal{S}}({\mathcal{EX}}({\mathcal{R}}))\subseteq[n]\times[m]$. ∎
@@ 4.4. Standard monomial basis and Hilbert series
Let $k$ be the number of shadow lines of $w$ that we use to construct ${\mathcal{S}}(w)$. Write $w\mapsto(P,Q)$ according to the Schensted correspondence, where $P,Q$ have the same shape $\lambda$. The shadow line construction in Subsection 4.1 implies that $k=\lambda_{1}$. In addition, Schensted correspondence possesses the standard property [Schensted_1961, Thm. 1] $\lambda_{1}=\mathrm{lis}(w)$, so we have $k=\mathrm{lis}(w)$ and thus Equation (4.4) holds. ∎
Surprisingly, although ${\mathcal{UZ}}_{n,m,r}$ looks larger than ${\mathfrak{S}}_{n}$, our extension operation $\mathcal{EX}$} embeds ${\mathcal{UZ}}_{n,m,r}$ into ${\mathfrak{S}}_{n+m-r}$, so that we can substitute n=Na,m=Nb,r=Nabn=N-a,m=N-b,r=N-a-b and hence embed ${\mathcal{UZ}}_{n,m,r}$ into ${\mathfrak{S}}_{N}$. Therefore, while [rhoades2024increasing, Corollary 3.13] states that
(4.5) ${\mathrm{Hilb}}({\mathbb{C}}[{\mathbf{x}}_{N\times N}]/I_{N,N,N};q)=\sum_{w\in{\mathfrak{S}}_{N}}q^{N-\mathrm{lis}(w)},$
the following modified version of Corollary 4.15 refines the index set of Equation (4.5).
Corollary 4.16.
For $N\in{\mathbb{Z}}_{>0}$ and $0\leq a,b<N$, we have
${\mathrm{Hilb}}({\mathbb{C}}[{\mathbf{x}}_{(N-a)\times(N-b)}]/I_{N-a,N-b,N-a-b};q)=\sum_{w}q^{N-\mathrm{lis}(w)}$
where $w$ ranges over $\{w\in{\mathfrak{S}}_{N}\,:\,\text{$w(1)<w(2)<\dots<w(a)$ and $w^{-1}(1)<w^{-1}(2)<\dots<w^{-1}(b)$}\}$.
Proof.
Write $\mathrm{Inc\coloneqq}\{w\in{\mathfrak{S}}_{N}\,:\,\text{$w(1)<w(2)<\dots<w(a)$ and $w^{-1}(1)<w^{-1}(2)<\dots<w^{-1}(b)$}\}$ and let n=Na,m=Nb,r=Nabn=N-a,m=N-b,r=N-a-b, then we have
$\mathrm{Inc}=\{w\in{\mathfrak{S}}_{n+m-r}\,:\,\text{$w(1)<w(2)<\dots<w(m-r)$ and $w^{-1}(1)<w^{-1}(2)<\dots<w^{-1}(n-r)$}\}.$
According to Corollary 4.15, it suffices to show that
${\mathcal{EX}}\,:\,{\mathcal{UZ}}_{n,m,r}$ $\longrightarrow\mathrm{Inc}$
${\mathcal{R}}$ $\longmapsto{\mathcal{EX}}({\mathcal{R}})$
is a bijection.
Since the diagram of $\mathcal{R}$} is contained in the diagram of ${\mathcal{EX}}({\mathcal{R}})$, it immediately follows that $\mathcal{EX}$} is injective. It remains to show that $\mathcal{EX}$} is surjective. For any $w\in\mathrm{Inc}$, choose maximal indices $1\leq i_{0}\leq m-r,1\leq j_{0}\leq n-r$ such that $w(i_{0})\leq n-r$ and $w^{-1}(j_{0})\leq m-r$. If such $i_{0}$ (resp. $j_{0}$) does not exist, let $i_{0}=0$ (resp. $j_{0}=0$). Then, we have $i_{0}]\cap w^{-1}(\{j_{0}+1,\dots,n-r\})=\varnothing$ and hence $w(i_{0})\leq j_{0}$. Similarly, we have $j_{0}]\cap w(\{i_{0}+1,\dots,m-r\})=\varnothing$ and hence $j_{0}\leq w(i_{0})$. These two inequalities force $w(i_{0})=j_{0}$. Therefore, $w(1)<w(2)<\dots<w(i_{0})$ indicates that $i_{0}\leq j_{0}$, while $w^{-1}(1)<w^{-1}(2)<\dots<w^{-1}(j_{0})$ indicates that $j_{0}\leq i_{0}$. As a result, we have $i_{0}=j_{0}$ and $w(i)=i$ for all $i\in[i_{0}]$, so $w={\mathcal{EX}}({\mathcal{R}})$ where ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r}$ arises from the removal of the first mrm-r columns and the first nrn-r rows from the diagram of $w$. ∎
@@ 6. Conclusion
We provide an open problem regarding log-concavity. Recall that a sequence of positive real numbers $(a_{1},a_{2},\dots,a_{N})$ is log-concave if for all $1<i<N$ we have $a_{i}^{2}\geq a_{i-1}\cdot a_{i+1}$. Chen conjectured [chen2008logconcavityqlogconvexityconjectureslongest] that the sequence $(a_{n,1},a_{n,2},\dots,a_{n,n})$ is log-concave where $a_{n,k}$ is the number of permutations in ${\mathfrak{S}}_{n}$ with longest subsequence of length $k$. Using the terminology of orbit harmonics, Chen’s conjecture means that the Hilbert series of the orbit harmonics ring $R({\mathfrak{S}}_{n})$ of the permutation matrix locus ${\mathfrak{S}}_{n}$ is log-concave.
We can generalize the definition of log-concavity by incorporating group action. Let $G$ be a group and $(V_{1},V_{2},\dots,V_{N})$ is a sequence of $G$-modules. We say that $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave if for all $1<i<N$ we have a $G$-module surjection $V_{i}\otimes V_{i}\twoheadrightarrow V_{i-1}\otimes V_{i+1}$ where the $G$-action is diagonal, i.e., $g\cdot(v\otimes w)=(g\cdot v)\otimes(g\cdot w)$. A graded $G$-module $V=\bigoplus_{d=0}^{N}V_{d}$ is $G$-equivariant log-concave if the sequence $(V_{1},V_{2},\dots,V_{N})$ is $G$-equivariant log-concave. Rhoades conjectured [rhoades2024increasing] that $R({\mathfrak{S}}_{n})$ is ${\mathfrak{S}}_{n}\times{\mathfrak{S}}_{n}$-equivariant log-concave, which generalizes Chen’s conjecture.
@@
Another interesting future direction is to study the orbit harmonics ring $R({\mathcal{Z}}_{n,m,r})$. Since ${\mathcal{Z}}_{n,m,r}\subseteq{\mathcal{UZ}}_{n,m,r}$, we need to add some elements to the generating set in Definition 3.1 to obtain the defining ideal ${\mathrm{gr}}\,{\mathbf{I}}({\mathcal{Z}}_{n,m,r})$, such as $\sum_{i=1}^{n}\sum_{j=1}^{m}x_{i,j}$ and ${\mathfrak{m}}({\mathcal{R}})$ for ${\mathcal{R}}\in{\mathcal{UZ}}_{n,m,r+1}$. Some relevant results are presented in Zhu [zhu2025rookplacementsorbitharmonics] focusing on the graded module structure.